OpenStudy (czarluc):

What is 3^9x when 27^x=4 And solve for (2^x)3x+3x=1 (two different questions^)

OpenStudy (anonymous):

\[27^x=4\] so \[3^{3x}=4\]

OpenStudy (czarluc):

but the bases are not the same :(

OpenStudy (anonymous):

they are now

OpenStudy (anonymous):

\[27=3^3\]so \[27^x=3^{3x}\]

OpenStudy (czarluc):

No, i mean 3 is not the same with 4

OpenStudy (anonymous):

since \[3^{3x}=4\] you can compute \[3^{9x}=(3^{3x})^3\]

OpenStudy (czarluc):

Yeah, but what about 4?

OpenStudy (anonymous):

the 4 is what you are going to cube

OpenStudy (anonymous):

\[3^{9x}=(3^{3x})^3=4^3\]

OpenStudy (czarluc):

What happens next?

OpenStudy (anonymous):

cube 4

OpenStudy (czarluc):

Why?

OpenStudy (anonymous):

nvm

OpenStudy (czarluc):

okok

OpenStudy (czarluc):

@satellite73 ?